# Galactica: A Scientific Large Language Model in Geoscience

### v1

@@ 6.2 Human Evaluation
ChatGPT is an AI language model developed by OpenAI, known for its ability to generate human-like text based on prompts, facilitate engaging conversations, answer questions, and perform a wide range of language-related tasks. 252525We use the 2023 March version of ChatGPT.
@@
For open-ended questions, the general large-scale model uses the interface output provided by ChatALL 262626https://github.com/sunner/ChatALL for consistency. Our large model interacts through our UI interface, where higher metric scores are preferred.
@@ Acknowledgement
This project is conducted under the DDE initiative. The computation resource was supported by the Advanced Computing East China Sub-center. This work is supported by NSF China (No.62020106005, 61960206002, 42050105, 62061146002, 62106143), National Key Technologies R&D Program (No. 2022YFB3904201), Shanghai Pilot Program for Basic Research - Shanghai Jiao Tong University. The second author would like to thank Wu Wen Jun Honorary Doctoral Scholarship, AI Institute, Shanghai Jiao Tong University.
@@ Appendix D Appendix: Prompts
ddescholar.abstract.title
@@
ddescholar.abstract.keyword
@@
ddescholar.reference.resolution
@@ Appendix F Appendix: Model Card
Further pre-train: A geoscience-related text corpus containing 65 billion tokens curated from extensive data sources in the big science project Deep-time Digital Earth (DDE), preserving as the largest geoscience-specific text corpus.
@@ H.1 Noun Definition
We present an example with prompt ”What is physical weathering and chemical weathering?“ as follows:
@@ H.2 Beginner Level Q&A
We present an example with prompt ”How does a cloud fill up with water?“ as follows:
@@ H.3 Intermediate Level Q&A
We present an example with prompt ”How does the movement of tectonic plates contribute to the formation of earthquakes and volcanic activity?“ as follows:
@@ H.4 Advanced Level Q&A
We present an example with prompt ”How did Earth and other planets form? Were planets formed in situ?“ as follows:
@@ H.5 Knowledge-based associative judgment question
We present an example with prompt ”What are the subfields that fall under magnetic polarity stratigraphy?“ as follows:
@@ H.7 Geoscience Research Functionality
We present an example with instruction ”Can you help me summarize the article?“ over the paper ”A challenge to the concept of slip-lines in extrusion tectonics“ as follows:
@@ Appendix I Appendix: Tool Learning Use cases
Example 2 Question: Introduce some arxiv paper about carbonate rock to me.
@@
• Earthquake is a way for the earth to release energy. Earthquake will produce ground vibration, strong vibration will cause reservoir dam break, mountain collapse, surface rupture, sand liquefaction, soft soil earthquake subsidence and ground settlement, resulting in earthquake disaster. Earthquake disaster has the characteristics of strong suddenness, great destructiveness, wide disaster scope, deep social impact, difficult defense, serious secondary disasters and so on, which is the first group disaster. For cities, due to the characteristics of dense population, numerous buildings, extensive infrastructure, concentrated wealth and multiple functions, once an earthquake occurs, it often produces huge disasters and extremely heavy losses. The article discusses the causes, effects, and consequences of earthquakes, as well as the damage they can cause to cities. Earthquakes are caused by the release of energy in the Earth’s crust and can result in a range of damages, including the collapse of mountains, the breaking of reservoir dams, ground surface ruptures, liquefaction of sand, subsidence of soft soils and settlements, all of which can lead to significant loss of life and property. Cities, with their high population densities, large numbers of buildings and infrastructure and diverse economic activities, are particularly vulnerable to damage from earthquakes. • The exploitation and utilization of groundwater resources meet the diversified needs of human beings for water resources, and achieve huge social and economic benefits. But at the same time, the development and utilization of groundwater resources also have a certain impact on the ecological environment, land salinization and desertification are serious, water pollution is serious, and a variety of negative environmental effects appear, seriously restricting the ecological sustainable development. In order to meet people’s water needs, groundwater has been over-exploited in the past time, which has caused frequent geological environmental problems, and has begun to affect people’s living environment. In this paper, the impact of groundwater exploitation on environmental geology is discussed, and the countermeasures are analyzed. This scientific article discusses the importance and benefits of using groundwater as a source of water for humans, but also acknowledges the negative impacts that can result from its overuse. The article highlights the consequences of overusing groundwater, such as land degradation, pollution and ecological damage, all of which can negatively impact the environment and human life. It is important to balance the need for groundwater with its potential negative effects, to ensure that it is used in a sustainable and responsible manner • Genetic engineering, as an important branch of biotechnology, has made remarkable progress. By precisely editing and modifying an organism’s genome, we open up entirely new possibilities for changing lives. This article will delve into the new chapter of genetic engineering, explore its application prospects in medicine, agriculture, environmental protection and other fields, and show its far-reaching impact on human society. This scientific article discusses the use and potential of Genetic Engineering in the field of agriculture. It highlights the ability of this technology to precisely edit and modify the genomes of organisms, allowing for the development of new products and methods that have the potential to greatly impact human life. The article explores the application of these techniques in fields such as medicine and agriculture and discusses their potential benefits and drawbacks. Overall, the article emphasizes the
@@
During the development and debugging phase, we encountered specific issues with the initialization of the Megatron Optimizer. It was observed that in the second step, the loss would increase to the level of randomly initialized models. We eliminated the possibility of errors in parameter conversion, as the same phenomenon persisted even with the use of Megatron’s openly released models for further pre-training. However, when we initialized the model from scratch, trained it for several steps, saved a checkpoint, and then loaded this checkpoint to resume training, the behavior normalized. We noted that the main difference between further pre-training and resuming training lies in the optimizer: in further pre-training, the optimizer is re-initialized, whereas in resuming training, it is imported from the checkpoint. This led us to hypothesize that there might be an issue with the optimizer’s initialization in the original code. Finally, we referred to a pull request 272727https://github.com/NVIDIA/Megatron-LM/pull/240/ on Github, which helped us identify and resolve this problem.
@@
Our primary training cluster is based on ROCm chips. During the development and debugging phase, we encountered another issue related to bugs in operators on ROCm. Initially, our tests on the ROCm cluster showed that training in FP16 usually worked well, producing relatively normal results. However, the outcomes for BF16 and FP32 were inaccurate, characterized by enormous gradients and losses. Interestingly, we couldn’t replicate this phenomenon in our local CUDA cluster, which operated using the same code, hyper-parameters, and checkpoints. Despite our numerous attempts, we found no solution until we came across a Github report 282828https://github.com/microsoft/Megatron-DeepSpeed/pull/96 that led us to identify and resolve the issue: a bug in the compilation of the LayerNorm operator’s source code in the ROCm environment. This discovery helped explain the observed behavior: FP16, with its smaller range of numerical values compared to BF16 and FP32, was less likely to exhibit such large gradients or losses, even though the accuracy in all these cases was affected by the bug.
Selection of Training Accuracy on ROCm
There are three types of training accuracies: FP16, BF16, and FP32. FP16, though relatively unstable, often grapples with numerical overflow. FP32 offers the best accuracy and stability but is the least efficient. BF16 strikes a balance between stability and efficiency, but its support is limited to a few devices. Initially, we preferred using BF16 in our ROCm cluster because FP16 presented challenges, as shown in our later experiments and evidenced in the chronicles of OPT training, and FP32 was significantly slower and rarely used in Large Language Model (LLM) training. However, our ROCm chips were not fully compatible with BF16; its BF16 computations largely depended on FP32 processing units in the hardware layer, lacking dedicated processing units, which meant no significant speedup over FP32. Considering these factors, we chose to use FP32 for further pre-training and supervised fine-tuning.

### v2

@@ 6.2 Human Evaluation
ChatGPT is an AI language model developed by OpenAI, known for its ability to generate human-like text based on prompts, facilitate engaging conversations, answer questions, and perform a wide range of language-related tasks. 282828We use the 2023 March version of ChatGPT.
@@
For open-ended questions, the general large-scale model uses the interface output provided by ChatALL 292929https://github.com/sunner/ChatALL for consistency. Our large model interacts through our UI interface, where higher metric scores are preferred.
@@ Acknowledgement
The computation resource was supported by the Advanced Computing East China Sub-center. This work is supported by NSF China (No.62020106005, 61960206002, 42050105, 62061146002, 62106143), National Key Technologies R&D Program (No. 2022YFB3904201), Shanghai Pilot Program for Basic Research - Shanghai Jiao Tong University. The second author would like to thank Wu Wen Jun Honorary Doctoral Scholarship, AI Institute, Shanghai Jiao Tong University.
@@ Appendix D Appendix: Prompts
deepliterature.abstract.title
@@
deepliterature.abstract.keyword
@@
deepliterature.reference.resolution
@@ Appendix F Appendix: Model Card
Further pre-train: A geoscience-related text corpus containing 65 billion tokens, preserving as the largest geoscience-specific text corpus.
@@ H.1 Noun Definition
We present an example with prompt ”What is physical weathering and chemical weathering¿‘ as follows:
[linecolor=baby_blue, linewidth=2pt, roundcorner=10pt, fontcolor=black, shadow=true, shadowsize=2pt, shadowcolor=light_grey]
ChatGPT:
Physical weathering and chemical weathering are two processes that contribute to the breakdown and alteration of rocks and minerals on Earth’s surface. Let’s explore each of them:
1. Physical Weathering: Physical weathering, also known as mechanical weathering, refers to the disintegration of rocks and minerals into smaller fragments without changing their chemical composition. It is primarily caused by physical forces and environmental factors.
2. Chemical Weathering: Chemical weathering involves the alteration or decomposition of rocks and minerals through chemical reactions. It occurs when minerals in rocks come into contact with water, atmospheric gases, or other substances that cause them to dissolve, decompose, or undergo chemical changes.
It’s important to note that physical and chemical weathering often occur simultaneously and can complement each other, accelerating the overall weathering process. Additionally, the type and rate of weathering are influenced by factors such as climate, rock composition, topography, and the presence of vegetation or organisms.
GeoGalactica
Physical weathering and chemical weathering are two of the main processes that cause rocks and minerals to break down and change over time.
Physical weathering refers to the breakdown of rocks and minerals due to the action of natural forces such as wind, water, ice, and gravity. These forces can cause rocks to crack, crumble, and disintegrate into smaller pieces. For example, when water flows over the surface of a rock, it can erode and wear away the rock’s surface, causing it to break apart.
Chemical weathering, on the other hand, refers to the breakdown of rocks and minerals due to the action of chemical reactions. These reactions occur when water, air, and other chemicals interact with the rocks and minerals, causing them to dissolve, decompose, or react with each other. For example, when carbon dioxide in the air dissolves in water, it forms carbonic acid, which can dissolve calcium carbonate in limestone, causing it to break down into calcium and bicarbonate ions.
In summary, physical weathering refers to the breakdown of rocks and minerals due to the action of natural forces, while chemical weathering refers to the breakdown of rocks and minerals due to the action of chemical reactions. Both processes play an important role in shaping the Earth’s landscape over time.
@@ H.2 Beginner Level Q&A
We present an example with prompt ”How does a cloud fill up with water¿‘ as follows:
[linecolor=baby_blue, linewidth=2pt, roundcorner=10pt, fontcolor=black, shadow=true, shadowsize=2pt, shadowcolor=light_grey]
ChatGPT:
"A cloud forms and ""fills up"" with water through a process known as condensation. It involves the transformation of water vapor into liquid water droplets or ice crystals, which are the visible components of clouds.
Here’s a simplified explanation of the process:
1. Evaporation: The sun’s heat causes water from various sources like oceans, lakes, and rivers to evaporate, turning it into water vapor. This water vapor is essentially invisible.
2. Moisture in the air: As the water evaporates, it rises into the atmosphere and mixes with the surrounding air. The capacity of the air to hold water vapor depends on factors such as temperature and humidity.
3. Cooling and saturation: As the warm, moist air rises higher into the atmosphere, it encounters cooler temperatures. Cooler air cannot hold as much moisture as warm air, so the water vapor begins to condense into tiny water droplets or ice crystals around microscopic particles in the air, such as dust, salt, or pollution. This process is called saturation.
4. Cloud formation: The condensed water droplets or ice crystals cluster together to form visible clouds. Clouds consist of countless tiny droplets or ice crystals suspended in the air, which give them their characteristic appearance.
The process of cloud formation is influenced by various factors such as air temperature, humidity levels, and the presence of condensation nuclei (particles that water vapor can condense onto). Different types of clouds, such as cumulus, stratus, or cirrus, form under specific atmospheric conditions and have distinct appearances.
It’s important to note that while clouds contain water, they don’t ""fill up"" like a container. The amount of water vapor in the atmosphere continuously changes as a result of various meteorological processes, including evaporation, condensation, and precipitation. Clouds form, evolve, and dissipate as these processes occur within the atmosphere."
GeoGalactica
A cloud fills up with water when water vapor in the air cools and condenses into tiny droplets or ice crystals. This happens when the temperature of the air drops below the dew point or frost point, which is the temperature at which the air can hold the maximum amount of water vapor.
When the temperature drops, the water molecules in the air slow down and collide with each other more frequently. This causes the water vapor to condense into tiny droplets or ice crystals, forming clouds. The more water vapor in the air, the more likely it is that the temperature will drop low enough for condensation to occur.
Clouds can also form when warm, moist air rises and cools. As the air rises, it expands and cools, causing the water vapor in the air to condense into tiny droplets or ice crystals. This process is known as adiabatic cooling.
In summary, a cloud fills up with water when water vapor in the air cools and condenses into tiny droplets or ice crystals. This can happen when the temperature of the air drops below the dew point or frost point, or when warm, moist air rises and cools.
@@ H.3 Intermediate Level Q&A
We present an example with prompt ”How does the movement of tectonic plates contribute to the formation of earthquakes and volcanic activity¿‘ as follows:
[linecolor=baby_blue, linewidth=2pt, roundcorner=10pt, fontcolor=black, shadow=true, shadowsize=2pt, shadowcolor=light_grey]
ChatGPT:
"The movement of tectonic plates is a key factor in the formation of earthquakes and volcanic activity. The Earth’s lithosphere, which consists of several large and numerous smaller tectonic plates, is divided into rigid pieces that float and move on the underlying semi-fluid asthenosphere. There are three main types of plate boundaries where interactions between these plates occur: divergent boundaries, convergent boundaries, and transform boundaries. At divergent boundaries, tectonic plates move away from each other, creating a gap or rift. This movement results in the upwelling of magma from the mantle, forming new crust. As the magma rises to the surface, it can lead to volcanic activity, with molten rock erupting to form volcanoes. However, the volcanic activity at divergent boundaries is generally less explosive compared to other types of volcanic activity. At convergent boundaries, tectonic plates collide with each other. There are three types of convergent boundaries: oceanic-oceanic, oceanic-continental, and continental-continental. When an oceanic plate converges with another oceanic plate or a continental plate, the denser oceanic plate is usually subducted (forced beneath) into the mantle. The subduction process can generate intense pressure and friction, resulting in the release of seismic energy and causing earthquakes. Additionally, as the subducted plate sinks into the mantle, it can melt, generating magma that rises to the surface and forms volcanoes. These volcanoes can be explosive and often create volcanic arcs, such as the Pacific Ring of Fire. At transform boundaries, tectonic plates slide past each other horizontally. The friction between the plates can cause them to become locked, preventing smooth movement. As the plates continue to push against each other, stress builds up along the locked boundary. Eventually, the stress overcomes the friction, leading to a sudden release of energy, resulting in earthquakes. Transform boundaries are primarily characterized by seismic activity rather than volcanic activity. In summary, the movement of tectonic plates at divergent boundaries leads to volcanic activity, while at convergent boundaries, it results in both earthquakes and volcanic activity. Transform boundaries primarily contribute to seismic activity. These interactions between tectonic plates play a fundamental role in shaping Earth’s surface and are responsible for the occurrence of earthquakes and the formation of volcanic features around the world."
GeoGalactica
"The movement of tectonic plates is the driving force behind the formation of earthquakes and volcanic activity on Earth. Tectonic plates are large, rigid sections of the Earth’s crust that move slowly over the surface of the planet. The movement of these plates is caused by convection currents in the Earth’s mantle, which is the layer beneath the crust and above the core.
When two tectonic plates move towards each other, they may collide, causing one plate to be forced under the other. This process is known as subduction, and it is responsible for the formation of many of the world’s mountain ranges, such as the Himalayas and the Andes. When two plates collide, the pressure can build up until it is released in a sudden, violent event known as an earthquake.
On the other hand, when two plates move in opposite directions, one may be forced upwards, causing the Earth’s crust to stretch and thin. This process is known as rifting, and it is responsible for the formation of many of the world’s rift valleys, such as the Great Rift Valley in Africa. When the crust is stretched too far, it can break apart, leading to the formation of new ocean basins.
Volcanic activity is also closely linked to the movement of tectonic plates. When two plates move towards each other, the pressure can build up until it is released in a sudden, violent event known as an earthquake. This can cause the Earth’s crust to crack."
@@ H.4 Advanced Level Q&A
We present an example with prompt ”How did Earth and other planets form? Were planets formed in situ¿‘ as follows:
[linecolor=baby_blue, linewidth=2pt, roundcorner=10pt, fontcolor=black, shadow=true, shadowsize=2pt, shadowcolor=light_grey]
ChatGPT:
"The formation of planets depends on several factors, including the gravitational pull of nearby celestial bodies, the temperature and pressure of the surrounding environment, and the availability of necessary resources for planet formation. Planets may form in situ (i.e., within their own gravitational field) or through the interaction with other celestial bodies. Some planets may form from dust particles in space, while others may form around existing celestial bodies such as comets or asteroids. The process of planet formation involves the accumulation of material over time until it reaches a critical mass where it can no longer be held back by gravity. This critical mass determines the final size and composition of the planet. Overall, the process of planet formation is complex and influenced by various factors."
GeoGalactica
"The formation of Earth and other planets is a complex process that is still not fully understood. However, there are several theories and models that attempt to explain the formation of our solar system and its planets.
One of the most widely accepted theories is the ""nebular hypothesis,"" which suggests that the solar system formed from a rotating cloud of gas and dust called the solar nebula. This cloud collapsed under its own gravity, forming the sun at its center. As the cloud collapsed, it spun faster and faster, forming a disk of gas and dust around the sun. This disk is what we now call the solar system, and it is from this disk that the planets formed.
One of the main challenges in understanding the formation of the solar system and its planets is the wide range of sizes and compositions of the planets. For example, Earth is a rocky planet with a solid surface, while Jupiter is a gas giant with a thick atmosphere. Scientists have developed several models to explain the formation of these different types of planets.
One of the most popular models for the formation of rocky planets like Earth is the ""core accretion model."" This model suggests that the first step in the formation of a rocky planet is the formation of a solid core. This core is made up of small solid particles, such as dust and rocks, that collide and stick together. As the core grows larger, it attracts more gas and dust from the solar nebula, forming a larger planet."
@@ H.5 Knowledge-based associative judgment question
We present an example with prompt ”What are the subfields that fall under magnetic polarity stratigraphy¿‘ as follows:
[linecolor=baby_blue, linewidth=2pt, roundcorner=10pt, fontcolor=black, shadow=true, shadowsize=2pt, shadowcolor=light_grey]
ChatGPT:
"There are four main subfields that fall under magnetic polarity stratigraphy: Stratigraphic magnetic polarity (SMP): SMP looks at the relative orientation of rocks and sediments along a given layer. It determines the direction of flow of the rocks and sediments in that layer. Tectonic magnetic polarity (TMP): TMP looks at the relationship between rocks and sediments in relation to the tectonic plate boundaries. It determines the relative movements of the plates and the rocks and sediments on either side. Paleomagnetism (PM): PM looks at the magnetic susceptibility of rocks and sediments. It determines the type of rock or sediment and the strength of the magnetic signal. Applied geomagnetism (AG): AG looks at the application of magnetic principles in different fields, such as geology, petrography, and mineralogy. It applies the principles of electromagnetic induction to solve problems in various fields. Overall, magnetic polarity stratigraphy falls under three main subfields: stratigraphic magnetic polarity (SMP), tectonic magnetic polarity (TMP), and applied geomagnetism (AG). These subfields combine to provide a comprehensive approach to analyzing the magnetic properties of rocks and sediments, and to interpreting their relationships with the underlying tectonic plate boundaries."
GeoGalactica
Magnetic polarity stratigraphy, also known as magnetostratigraphy, is a subfield of geophysics and geology that uses the polarity of the Earth’s magnetic field to date and correlate sedimentary rock layers. The Earth’s magnetic field is generated by the movement of molten iron in the Earth’s outer core, and the polarity of the magnetic field can be either positive or negative. The polarity of the Earth’s magnetic field has changed over time, and these changes can be used to correlate sedimentary rock layers from one location to another.
There are several subfields that fall under magnetic polarity stratigraphy, including:
1. Magnetostratigraphy: This subfield uses the polarity of the Earth’s magnetic field to date and correlate sedimentary rock layers.
2. Biostratigraphy: This subfield uses the distribution of fossils in sedimentary rock layers to date and correlate sedimentary rock layers.
3. Lithostratigraphy: This subfield uses the physical characteristics of sedimentary rock layers to date and correlate sedimentary rock layers.
4. Chemostratigraphy: This subfield uses the chemical composition of sedimentary rock layers to date and correlate sedimentary rock layers.
5. Magneto-biostratigraphy: This subfield combines magnetostratigraphy and biostratigraphy to date and correlate sedimentary rock layers.
6. Magneto-lithostratigraphy: This subfield combines magnetostratigraphy and lithostratigraphy to date and correlate sedimentary rock layers.
@@ H.6 Research Paper Titling Task
[linecolor=baby_blue, linewidth=2pt, roundcorner=10pt, fontcolor=black, shadow=true, shadowsize=2pt, shadowcolor=light_grey]
Original title:
Analysis of the geo-hazards triggered by the 12 May 2008 Wenchuan Earthquake, China
ChatGPT:
Spatial Distribution and Characteristics of Geo-Hazards Triggered by the Wenchuan Earthquake: Focus on Landslides
GeoGalactica
Analysis of the Distribution of Geo-Hazards Triggered by the 2008 Wenchuan Earthquake
@@ H.7 Geoscience Research Functionality
We present an example with instruction ”Can you help me summarize the article¿‘ over the paper ”A challenge to the concept of slip-lines in extrusion tectonics“ as follows:
[linecolor=baby_blue, linewidth=2pt, roundcorner=10pt, fontcolor=black, shadow=true, shadowsize=2pt, shadowcolor=light_grey]
ChatGPT:
The article discusses wide-open V-shaped conjugate strike-slip faults in Asia and their association with extrusion tectonics. It highlights critical problems with the tectonic model based on the slip-line theory of plasticity. According to the theory, the angles between conjugate slip-lines should be normal to each other, but in reality, they are always greater than 90 degrees (on average around 110 degrees) in the direction of contraction.
Another issue is the failure of the slip-line theory to explain cases where the extrusion rate is significantly higher than the indent rate, as observed in the Anatolian area in the eastern Mediterranean. The article proposes that these problems can be addressed by considering the Maximum-Effective-Moment (MEM) Criterion. The MEM Criterion predicts that shear zones’ orientations should theoretically be at an angle of 54.7 degrees (or practically around 55 degrees with the contractional direction), which fundamentally controls the orientations of the strike-slip faults involved in extrusion tectonics.
The concept of escape or extrusion tectonics, where India acts as an indentor and Tibet and neighboring areas act as a plastic body, is widely accepted to explain the tectonic patterns in various regions. However, the slip-line theory on which this model is based is found to be incompatible with field observations and experiments. The angles between conjugate ductile shear zones, known as ""V-shaped conjugate faults,"" are consistently obtuse in nature, contrary to the slip-line theory’s predictions.
These wide-open V-shaped conjugate strike-slip faults are prevalent in central Tibet, the eastern Alps, western Turkey, eastern Afghanistan, western Mongolia, Southeast Asia, and the Gulf of Thailand. They typically lie at an angle of 60 to 75 degrees from the maximum compressive stress direction, with a conjugate angle ranging from 120 to 145 degrees in the s1 direction. Due to their obtuse angles, they can be mistakenly identified as normal conjugate faults with acute angles in the s1-direction. The article suggests using the term ""wide-open"" to distinguish this type of conjugate fault.
In summary, the article focuses on the wide conjugate angles observed in wide-open V-shaped conjugate strike-slip faults in Asia, which are associated with extrusion tectonics. It highlights the limitations of the slip-line theory and proposes the use of the MEM Criterion to explain the fault orientations. The article also discusses the prevalence of these faults in various regions and suggests adding the term ""wide-open"" to describe their distinct characteristics.
GeoGalactica
The passage discusses the issue of wide-open V-shaped conjugate strike-slip faults in Asia, which are typically related to extrusion tectonics. However, the tectonic model based on the slip-line theory of plasticity has critical problems associated with it. According to the theory of plasticity, the conjugate sets of slip-lines in plane deformation should be normal to each other, but in reality, the angles between the conjugate strike-slip faults in the eastern Mediterranean, Tibet-middle Asia, China, and the Indochina Peninsular regions are always wider than 90 degrees in the direction of contraction. Additionally, the slip-line theory fails to explain how, in some cases, such as in the Anatolian area in the eastern Mediterranean, the extrusion rate is much higher than the indent rate.
@@ Appendix I Appendix: Tool Learning Use cases
Example 2
Question: Introduce some arxiv paper about carbonate rock to me.
@@
[linecolor=baby_blue, linewidth=2pt, roundcorner=10pt, fontcolor=black, shadow=true, shadowsize=2pt, shadowcolor=light_grey, nobreak=true]
Please write a paper on the study of Dinosaur Fossils.
As an AI, I do not have my own personal opinion, but I can provide information and facts about the topic. Plate tectonics is the scientific theory that explains the movement of large sections of Earth’s crust, called plates. These plates are made of rigid, solid rock that move slowly over the surface of our planet. The movement is caused by the convection currents in the mantle, the layer of hot, molten rock beneath our crust. As the plates move, they can collide with each other, causing mountain ranges to form, or move apart, creating new ocean basins. This movement can also cause earthquakes, volcanic eruptions, and changes in climate. Overall, plate tectonics
Please write a paper on the Climate Change study.
Climate change is one of the most pressing environmental issues of our time. It is caused by the release of greenhouse gases, such as carbon dioxide, methane, and nitrous oxide, into the atmosphere. These gases trap heat from the sun and cause the Earth’s temperature to rise, leading to changes in weather patterns, rising sea levels, more frequent extreme weather events and loss of biodiversity. Climate change has far-reaching consequences for our planet and its inhabitants, making it essential that we understand its causes and effects. In this paper, we will explore the study of climate change, including its history, causes, effects and potential solutions. We will also look at the role that scientists and researchers play in understanding and addressing this global issue.
Please write a paper on the research of Urbanization and Social Change.
Urbanization is one of the most significant social, economic, and cultural transformations of our time. It is the process by which people move from rural areas to urban centers, resulting in the growth and expansion of cities and towns. This process has been ongoing for centuries, but the pace and scale of urbanization has accelerated dramatically in recent decades, driven by factors such as population growth, globalization, technological advancements and economic development. As a result, the world is now more urbanized than ever before, with more than half the global population now living in urban areas. However, urbanization is not without its challenges, as it is often associated with social inequality, poverty, environmental degradation and other forms of social change. In this paper, we will explore the various ways in which urbanization and social changes are intertwined, examining both the positive and negative impacts of this process on society. We will also discuss the role of government, NGOs and the private sector in addressing these challenges and promoting equitable and sustainable urbanization.
@@
[linecolor=baby_blue, linewidth=2pt, roundcorner=10pt, fontcolor=black, shadow=true, shadowsize=2pt, shadowcolor=light_grey]
Earth plate movement: The outer part of the Earth is divided into several large plates, which influence the shape and seismic activity of the crust in a way that is constantly moving and colliding.
As an AI, I do not have my own personal opinion, but I can provide information and facts about the topic. Plate tectonics is the scientific theory that explains the movement of large sections of Earth’s crust, called plates. These plates are made of rigid, solid rock that move slowly over the surface of our planet. The movement is caused by the convection currents in the mantle, the layer of hot, molten rock beneath our crust. As the plates move, they can collide with each other, causing mountain ranges to form, or move apart, creating new ocean basins. This movement can also cause earthquakes, volcanic eruptions, and changes in climate. Overall, plate tectonics…
Earth’s water cycle: The cycle of water in the Earth, including evaporation, condensation, precipitation, surface runoff, and groundwater.
As an AI, I don’t have my own personal opinion, but I can provide information and facts about the topic. The Earth’s Water Cycle is a natural process that involves the continuous movement and transformation of liquid water, water vapor and ice on, above and below the surface of the planet. Water evaporates from the ocean, lakes and land surfaces, then rises into the atmosphere where it cools and condenses into clouds. When the clouds become heavy with water droplets or ice crystals, they fall back to the ground as precipitation such as rain, snow, hail or sleet. Some of this precipitation flows into rivers, streams and lakes, while the rest recharges the groundwater or runs off into oceans. This cycle is essential for life on Earth as it helps regulate the climate, supports plant growth and sustains the water supply for humans, animals and other living organisms. However
Earth’s climate changes: The Earth’s climate changes over time, either naturally or as a result of human activities.
As an AI, I do not have my own personal opinions, but I can provide information and facts about the topic. Climate change refers to the long-term changes in weather patterns, temperature, and precipitation that occur on Earth. These changes are caused by a combination of natural factors, such as variations in solar activity and volcanic eruptions, as well as human activity, including the burning of fossil fuels and deforestation. According to scientific evidence, human-caused climate change is causing global temperatures to rise, leading to more frequent and intense heatwaves, droughts, hurricanes and other extreme weather events. It is important to reduce our carbon emissions and transition to renewable energy sources in order to slow down and potentially reverse these changes. However, it is essential that we continue to research and understand the complex interactions between the Earth and its climate, so we can make informed decisions and take action to protect
@@
[linecolor=baby_blue, linewidth=2pt, roundcorner=10pt, fontcolor=black, shadow=true, shadowsize=2pt, shadowcolor=light_grey]
Earthquake is a way for the earth to release energy. Earthquake will produce ground vibration, strong vibration will cause reservoir dam break, mountain collapse, surface rupture, sand liquefaction, soft soil earthquake subsidence and ground settlement, resulting in earthquake disaster. Earthquake disaster has the characteristics of strong suddenness, great destructiveness, wide disaster scope, deep social impact, difficult defense, serious secondary disasters and so on, which is the first group disaster. For cities, due to the characteristics of dense population, numerous buildings, extensive infrastructure, concentrated wealth and multiple functions, once an earthquake occurs, it often produces huge disasters and extremely heavy losses.
The article discusses the causes, effects, and consequences of earthquakes, as well as the damage they can cause to cities. Earthquakes are caused by the release of energy in the Earth’s crust and can result in a range of damages, including the collapse of mountains, the breaking of reservoir dams, ground surface ruptures, liquefaction of sand, subsidence of soft soils and settlements, all of which can lead to significant loss of life and property. Cities, with their high population densities, large numbers of buildings and infrastructure and diverse economic activities, are particularly vulnerable to damage from earthquakes.
The exploitation and utilization of groundwater resources meet the diversified needs of human beings for water resources, and achieve huge social and economic benefits. But at the same time, the development and utilization of groundwater resources also have a certain impact on the ecological environment, land salinization and desertification are serious, water pollution is serious, and a variety of negative environmental effects appear, seriously restricting the ecological sustainable development. In order to meet people’s water needs, groundwater has been over-exploited in the past time, which has caused frequent geological environmental problems, and has begun to affect people’s living environment. In this paper, the impact of groundwater exploitation on environmental geology is discussed, and the countermeasures are analyzed.
This scientific article discusses the importance and benefits of using groundwater as a source of water for humans, but also acknowledges the negative impacts that can result from its overuse. The article highlights the consequences of overusing groundwater, such as land degradation, pollution and ecological damage, all of which can negatively impact the environment and human life. It is important to balance the need for groundwater with its potential negative effects, to ensure that it is used in a sustainable and responsible manner
Genetic engineering, as an important branch of biotechnology, has made remarkable progress. By precisely editing and modifying an organism’s genome, we open up entirely new possibilities for changing lives. This article will delve into the new chapter of genetic engineering, explore its application prospects in medicine, agriculture, environmental protection and other fields, and show its far-reaching impact on human society.
This scientific article discusses the use and potential of Genetic Engineering in the field of agriculture. It highlights the ability of this technology to precisely edit and modify the genomes of organisms, allowing for the development of new products and methods that have the potential to greatly impact human life. The article explores the application of these techniques in fields such as medicine and agriculture and discusses their potential benefits and drawbacks. Overall, the article emphasizes the
@@
During the development and debugging phase, we encountered specific issues with the initialization of the Megatron Optimizer. It was observed that in the second step, the loss would increase to the level of randomly initialized models. We eliminated the possibility of errors in parameter conversion, as the same phenomenon persisted even with the use of Megatron’s openly released models for further pre-training. However, when we initialized the model from scratch, trained it for several steps, saved a checkpoint, and then loaded this checkpoint to resume training, the behavior normalized. We noted that the main difference between further pre-training and resuming training lies in the optimizer: in further pre-training, the optimizer is re-initialized, whereas in resuming training, it is imported from the checkpoint. This led us to hypothesize that there might be an issue with the optimizer’s initialization in the original code. Finally, we referred to a pull request 303030https://github.com/NVIDIA/Megatron-LM/pull/240/ on Github, which helped us identify and resolve this problem.
@@
Our primary training cluster is based on ROCm chips. During the development and debugging phase, we encountered another issue related to bugs in operators on ROCm. Initially, our tests on the ROCm cluster showed that training in FP16 usually worked well, producing relatively normal results. However, the outcomes for BF16 and FP32 were inaccurate, characterized by enormous gradients and losses. Interestingly, we couldn’t replicate this phenomenon in our local CUDA cluster, which operated using the same code, hyper-parameters, and checkpoints. Despite our numerous attempts, we found no solution until we came across a Github report 313131https://github.com/microsoft/Megatron-DeepSpeed/pull/96 that led us to identify and resolve the issue: a bug in the compilation of the LayerNorm operator’s source code in the ROCm environment. This discovery helped explain the observed behavior: FP16, with its smaller range of numerical values compared to BF16 and FP32, was less likely to exhibit such large gradients or losses, even though the precision in all these cases was affected by the bug.
Selection of Training Precision on ROCm
There are three types of training precisions: FP16, BF16, and FP32. FP16, though relatively unstable, often grapples with numerical overflow. FP32 offers the best precision and stability but is the least efficient. BF16 strikes a balance between stability and efficiency, but its support is limited to a few devices. Initially, we preferred using BF16 in our ROCm cluster because FP16 presented challenges, as shown in our later experiments and evidenced in the chronicles of OPT training, and FP32 was significantly slower and rarely used in Large Language Model (LLM) training. However, our ROCm chips were not fully compatible with BF16; its BF16 computations largely depended on FP32 processing units in the hardware layer, lacking dedicated processing units, which meant no significant speedup over FP32. Considering these factors, we chose to use FP32 for further pre-training and supervised fine-tuning.